Yuta, Manasa, Jamie, Jenne, Steve, Rana
Starting this morning, we removed the temporary half PRC mirror in front of BS and started to align the IFO in prep for an in-air lock of the PRMI.
This morning, using the new awesome steerable active input TTs, Jenne and I centred the beam on PRM, PR2/3, BS, ITMY and ETMY.
After lunch, Yuta and Manasa aligned the Y ARM, by looking at the multi-pass beam. The X-end door was still on, so they roughly aligned to the X ARM by centring on ITMX with BS. They then got fringes at the BS, and tweaked the ITMs and PRM to get full fringes at BS.
We're currently stuck because the REFL beam appears to be clipped coming out of the faraday, even though the retro-reflected beam from PRM is cleanly going through the faraday output aperture. The best guess at the moment is that the beam is leaving MC at an angle, so the retro-reflected beam is coming out of the faraday at an angle. We did not center spots on MC mirrors before we started the alignment procedure today. That was dumb.
We may be ok to do our PRMI characterization with the clipped REFL, though, then we can fix everything right before we close up. We're going to need to go back to touch up alignment before we close up anyway (we need to get PR2 centered).
Yuta and Manasa are finishing up now by making sure the AS and REFL beams are cleanly existing onto the AS table.
Tomorrow we will set up the PRM oplev, and start to look at the in-air PRMI. Hopefully we can be ready to close up by the end of the week.